If the perimeter of a rectangle is 24 ft and the width is 7 ft, find its length.
Found 2 solutions by fractalier, checkley71:
Answer by fractalier(2101)

Perimeter P = 2L + 2W
If we know P and W, plug in to find L...
24 = 2L + 2(7)
24 = 2L + 14
10 = 2L
L = 5 ft

Answer by checkley71(8403)

2*length+2*width=perimeter then 2L+2*7=24 OR 2L+14=24 OR 2L=24-14 OR 2L=10 OR
L=10/2 OR L=5